OpenStudy (anonymous):

Consider the functions f(x) = 2x + 1 and g(x) = x2 − 10. What is the value of (f − g)(5)? −4 4 −11 26

OpenStudy (amorfide):

just subtract the terms, then sub x=5

OpenStudy (unklerhaukus):

\[(f-g)(x)=f(x)-g(x)=(2x+1)-(x^2-10)\] \[(f-g)(5)=f(5)-g(5)=(2\times5+1)-(5^2-10)\]

OpenStudy (anonymous):

-4?

OpenStudy (unklerhaukus):

good show

OpenStudy (anonymous):

thxs for the help